# Modeling fast X-ray variability around an accreting black hole

### v1

@@ Abstract
Hard lags, characterized by high-energy (hard) photons lagging behind low-energy (soft) photons, are typically observed at low Fourier frequencies (a few hertz and below). These hard lags are thought to arise from the propagation of fluctuations in the mass accretion rate from the cold outer regions to the hot inner regions. As a result, soft photons respond more quickly than hard photons (Kotov et al., 2001; Ingram & Done, 2012). In all BHXRBs with sufficiently high quality observations, the lags can exhibit a reversal of sign at high frequencies, resulting in soft lags, which have the opposite definition to hard lags (Wang et al., 2022a; Marco & Ponti, 2019; Chand et al., 2022; Yu et al., 2023). Soft lags are usually explained by the time delay caused by the reflection process, i.e., reverberation mapping, where the disc reflects hard photons from the hot X-ray source. The reflected photons take a longer path to reach the observer compared to the directly emitted Comptonized photons (Emmanoulopoulos et al., 2014; Uttley et al., 2014). Recently, Uttley & Malzac(2024) proposed that the Comptonized photons would heat up the disc to enhance variable disc thermal emission, which could contribute to the soft lags. The zero-crossing point is defined as the frequency at which the soft lag transitions to hard lag. The zero-crossing point at high frequencies is thought to provide insights into the geometry of the inner accretion flow and its evolution (Mizumoto et al., 2018; Marco et al., 2021). However, there is still no consensus on the relationship between the zero-crossing points and the physical properties of the inner accretion flow.
@@
In recent years, the analysis of timing properties in BHXRBs has typically revealed the presence and evolution of hard and soft lags, such as in MAXI J1820+070, MAXI J1535-571 and GX 339–4 (De Marco et al., 2017; Kara et al., 2019; Wang et al., 2021; Yu et al., 2023). For example, in the outburst of MAXI J1820+070, hard lags are detected in low-frequency bands below several hertz, whereas soft lags dominate higher-frequency bands above ten hertz, with amplitudes ranging from approximately 0.2 to 3 milliseconds. During the rising hard state, a trend for soft lags towards a shorter timescale is observed, where their amplitudes are decreasing (Kara et al., 2019), alongside zero-crossing points shifting to higher frequencies (Marco et al., 2021).
@@
It is typically thought that hard lags and soft lags arise from viscous processes and reverberation delays (Ingram & Done, 2012; Ingram & van der Klis, 2013; Emmanoulopoulos et al., 2014). The evolution of these lags is believed to be driven by the inner accretion flow (Alston et al., 2020; Wang et al., 2022a; Kalemci et al., 2022). Some theoretical models, such as Reltrans (Ingram et al., 2019), provide explanations for the observed lag patterns. The Reltrans model assumes a simple lamppost geometry for the hard X-ray source, which illuminates the thin disc with a cutoff power-law spectrum. The "reflection" spectrum re-emitted from the disk features a prominent $\sim 6.4$ keV iron line, a $\sim 20-30$ keV Compton hump, and a strong soft excess. The light crossing delay between direct and reflected X-rays generates a soft lag between the flux in an energy band dominated by the reflection soft excess and the flux in a continuum-dominated band (García et al., 2013). However, observational studies imply that the X-ray source is typically radially extended (Krawczynski et al., 2022; Zhang et al., 2022; You et al., 2023; Wang & Zhang, 2024) and/or vertically extended (Kylafis et al., 2008; Ma et al., 2021; You et al., 2021; Méndez et al., 2022). Wilkins et al. (2016) provided a model that successfully reproduced the low-frequency hard lags and high-frequency soft lags with a corona both radially and vertically extended by assuming an energy descent in the corona and the bulk effect of the corona. Though this model explains the observed lags patterns with reverberation lags, spectral fitting indicates multiple emission components within the low-energy band (Marco et al., 2021; Wang et al., 2021; You et al., 2021), suggesting that inter-band lags may result from these multiple components. Thus, a quantitative simulation for frequency-dependent lag spectra that considers an extended corona and self-consistent radiative transfer is necessary. Furthermore, a comprehensive model is essential to explore the connection between the inner accretion flow and the evolution of the lags.
@@ 2.1 Geometry
Figure 1: Cross section of the hyperbolic corona floating above the BH. The upper and lower boundaries are controlled by the height parameters, $h_{t},h_{b}$. The outer surface is a hyperboloid with semi-major axis $a$ and semi-minor axis $b$
We consider a geometrically thin and optically thick disc that extends from the inner radius $r_{\text{in}}$ to the outer radius $r_{\text{out}}$. We model the corona as a hyperbolic surface with lower and upper surfaces (see Fig. 1). This geometry is most readily interpreted as the base of the jet. Whether this geometry is vertically or radially extended with respect to the disc depends on the parameters used. The hyperboloid is parameterized by the following equations:
@@
where $a$ and $b$ are the semi-major and semi-minor axis of the hyperboloid, respectively. $h_{t}$ and $h_{b}$ are the top and bottom boundaries in $z$-direction. $\{x,y,z\}$ are the Cartesian coordinates for the black hole frame. We mainly consider that the corona is wider than it is tall, with the parameters $a=10R_{\rm g},b=2R_{\rm g}$. An alternative spherical corona case will be discussed in Sect. 5. Some key parameters are shown in Table 1, with their definition and default values.
@@ 3 Monte Carlo Simulation
The code traces one single photon at one time. The trajectories of the photons are determined geometrically within the Newtonian framework for simplicity. This means light bending effect is not considered, which would result in a certain lack of reality. There would be an improvement in our future model.
@@
A fraction of photons from the disc will strike the corona. These photons may undergo Comptonization, which would contribute to the Compton component. For simplicity, this work does not consider the effects of outflow or inflow velocity, nor the precession of the corona. Additionally, it assumes a uniform plasma density and temperature within the corona. The simulation of Comptonization follows the standard methods established by Pozdnyakov et al.(1983) and Gorecki & Wilczewski(1984), which are parameterized by several factors related to the properties of the corona, such as optical depth and electron density, to determine whether the photons will be scattered.
@@ 3.1 Implementation of disc variability
In this study, we propose that the rapid variability originates from fluctuations in the accretion rate of the disc. To model the fluctuating accretion flow, we utilize the fluctuation model proposed by Ingram & Done(2011; 2012), which addresses local fluctuations and their propagation. The fluctuations propagate inward from the outer radius $r_{\text{out}}$ to the inner radius $r_{\text{in}}$ through the accretion flow, merging with local fluctuations along the way. For simplicity, we divide the accretion flow into $N_{\rm r}$ radial bins, each centered at $r_{n}$ in a logarithmic scale. The power spectrum of the local fluctuation at radius $r_{n}$ is represented by a zero-centered Lorentzian function:
@@
where $f$ is frequency, $f_{\mathrm{visc}}(r_{n})=f_{\max}(r_{n}/r_{i})^{-3/2}$ is the viscous frequency at radius $r_{n}$, in which $r_{i}$ is the inner edge of the disc (Shakura & Sunyaev, 1973; Rapisarda, 2017). $f_{\max}$ is a free model parameter that sets the maximum viscous frequency in the disc.
@@ 3.2 Implementation of disc reflection
Partial Comptonized photons from the corona are reflected by the disc, resulting in both emission line and continuum emission. In the works of You et al. (2018, 2020), the reflection for a neutral disc was simulated to study the QPO variation. However, previous studies on BHXRBs have suggested the presence of an ionized disc (You et al., 2021). It has been shown that the flux in the soft band, for instance, 0.1-1 keV, is highly dependent on the ionization state of the disc (García et al., 2013). Therefore, simulating the reflection of an ionized disc is essential. This involves modeling various processes such as absorption, excitation, and photoionization, which can be quite time-consuming. To address this, we adopt the table model Xillver (García et al., 2013) to simulate disc reflection.
Figure 2: Schematic of the two-zone method. The corona is separated into two zones according to an uncovering length $\lambda$, with transmittance $\kappa_{1}$ and $\kappa_{2}$, respectively. Denoting the radius of the vertical projection of the corona on the disc by $r_{\text{cover}}$ and the reflection emission $\mathcal{F}(r,E,t)$ at radius $r$, when $r<r_{\text{cover}}-\lambda$, $\mathcal{F}_{\text{refl}}(r,E,t)=\kappa_{1}\tilde{\mathcal{F}}_{\text{refl}}(r,E,t)$ and when $r_{\text{cover}}-\lambda<r<r_{\text{cover}}$, $\mathcal{F}_{\text{refl}}(r,E,t)=\kappa_{2}\tilde{\mathcal{F}}_{\text{refl}}(r,E,t)$.
@@
4 Result
We aim to simulate the inter-band lags across Fourier frequencies in the BHXRB. This will be achieved through Monte Carlo simulations of the rapid variability within the disc-corona system, taking into account both propagating fluctuations and reverberation mapping. Our goal is to quantitatively test whether these propagating fluctuations and reverberation mapping can explain the observed hard lag at low frequencies and the soft lag at high frequencies. Additionally, we will compare the results of our simulations with the observed evolution of frequency-dependent lags, which will help us constrain the geometry of the accretion flow and its evolution during outbursts. In this work, we will conduct simulations and comparisons specifically for MAXI J1820+070, which has provided valuable observations of fast variability (Kara et al., 2019; Marco et al., 2021; Kawamura et al., 2023).
[…]
@@ 4 Result
The time-averaged energy spectrum of the three components is presented in Fig. 3 for demonstration purposes. The spectrum exhibits a cutoff power-law component with photon index $\sim 2$, a thermal component with the peak temperature $\sim 0.1$ keV, and a high-ionized reflection component. These features are consistent with the spectral fitting results, e.g., You et al. (2021) and Gao et al. (2023). We focus on an energy band of interest ranging from 0.1 to 0.3 keV, where all components are comparable, and a reference band from 2 to 5 keV, where the Compton component is prominent. The lag spectra between these two bands can be calculated using the Python package Stingray (Huppenkothen et al., 2019), as shown in Fig. 4. The errors arise from the incoherent components between the two bands and the stochastic nature of the variability (Bendat & Piersol, 2011; Uttley et al., 2014). The solid line in the plot represents a linear interpolation to guide the eye. We observe a low-frequency hard lag and a high-frequency soft lag in the spectrum, which are consistent with observations(Kara et al., 2019; Yu et al., 2023; Marco et al., 2021).
@@
A trend towards shorter timescales for the lags is observed in MAXI J1820+070 as the accretion rate increases (Kara et al., 2019). Moreover, Marco et al. (2021) note that the movement of the zero-crossing points from low to high frequencies indicates an evolution of the physical properties of the inner accretion flow (Alston et al., 2020; Wang et al., 2022a; Kalemci et al., 2022). We attribute these physical properties to the geometry of the corona and the inner viscosity. Our simulations captured the evolution of the corona’s geometrical size and viscosity frequency, with the corona’s upper bound decreasing from $35~{}R_{\rm g}$ to $16~{}R_{\rm g}$, accompanied by an increase in the maximum viscous frequency from 2 Hz to 30 Hz. Four evolving states are simulated, labeled state-1 to state-4, where state-1 is characterized by a maximum viscous frequency of 2 Hz and an upper bound for the corona of $35~{}R_{\rm g}$. The lag spectra between 0.1-0.3 keV and 2-5 keV for the four states are presented in Fig. 5.
@@ 5 Discussion
In typical timing spectral analysis, the energy-dependent lag spectrum is a good probe to understand the physical processes behind the lags at different energies in a specific frequency band (Arévalo et al., 2006; Cassatella et al., 2012; Marco et al., 2015). The energy-dependent lag spectrum of MAXI J1820+070 shows that in the low frequency of 0.1-1 Hz, the lags increase as the energies increase in the energy band of 0.5-1 keV (Kara et al., 2019). This may originate from the propagation of fluctuations from a low energy area to high energy area, causing propagating lags. We tried to reproduce the energy-dependent lag spectrum from our simulation. We found that the simulated lag matches the the trend of the observed spectrum in the energy band below 1 keV, however, it breaks to a constant in the energy band over 1 keV. It’s necessary to note that the disk component vanishes over 1 keV, as shown in Fig. 3, and propagation of fluctuation in the corona is not considered in this work. Thus, after the disk component vanishes, the fluctuation can not propagate to the higher energy area, which may explain why the lags do not increase over 1 keV. We will implement the propagation of fluctuation in the corona in future work.

### v2

@@ Abstract
Hard lags, characterized by high-energy (hard) photons lagging behind low-energy (soft) photons, are typically observed at low Fourier frequencies (a few hertz and below). These hard lags are thought to arise from the propagation of fluctuations in the mass accretion rate from the cold outer regions to the hot inner regions. As a result, soft photons respond more quickly than hard photons (Kotov et al., 2001; Ingram & Done, 2012). In all BHXRBs with sufficiently high-quality observations, the lags can exhibit a reversal of sign at high frequencies, resulting in soft lags, which have the opposite definition to hard lags (Wang et al., 2022a; De Marco & Ponti, 2019; Chand et al., 2022; Yu et al., 2023). Soft lags are usually explained by the time delay caused by the reflection process, i.e., reverberation mapping, where the disc reflects hard photons from the hot X-ray source. The reflected photons take a longer path to reach the observer compared to the directly emitted Comptonized photons (Emmanoulopoulos et al., 2014; Uttley et al., 2014; Wilkins & Fabian, 2013). Recently, Uttley & Malzac (2024) proposed that the Comptonized photons would heat up the disc to enhance variable disc thermal emission, which could contribute to the soft lags. The zero-crossing point is defined as the frequency at which the soft lag transitions to hard lag. The zero-crossing point at high frequencies is thought to provide insights into the geometry of the inner accretion flow and its evolution (Mizumoto et al., 2018; De Marco et al., 2021). However, there is still no consensus on the relationship between the zero-crossing points and the physical properties of the inner accretion flow.
@@
In recent years, the analysis of timing properties in BHXRBs has typically revealed the presence and evolution of hard and soft lags, such as in MAXI J1820+070, MAXI J1535-571 and GX 339–4 (De Marco et al., 2017; Kara et al., 2019; Wang et al., 2021; Yu et al., 2023). For example, in the outburst of MAXI J1820+070, hard lags are detected in low-frequency bands below several hertz, whereas soft lags dominate higher-frequency bands above ten hertz, with amplitudes ranging from approximately 0.2 to 3 milliseconds. During the rising hard state, a trend for soft lags towards a shorter timescale is observed, where their amplitudes are decreasing (Kara et al., 2019), alongside zero-crossing points shifting to higher frequencies (De Marco et al., 2021).
@@
It is typically thought that hard lags and soft lags arise from viscous processes and reverberation delays (Ingram & Done, 2012; Ingram & van der Klis, 2013; Emmanoulopoulos et al., 2014). The evolution of these lags is believed to be driven by the inner accretion flow (Alston et al., 2020; Wang et al., 2022a; Kalemci et al., 2022). Some theoretical models, such as Reltrans (Ingram et al., 2019), provide explanations for the observed lag patterns. The Reltrans model assumes a simple lamppost geometry for the hard X-ray source, which illuminates the thin disc with a cutoff power-law spectrum. The "reflection" spectrum re-emitted from the disk features a prominent $\sim 6.4$ keV iron line, a $\sim 20-30$ keV Compton hump, and a strong soft excess. The light crossing delay between direct and reflected X-rays generates a soft lag between the flux in an energy band dominated by the reflection soft excess and the flux in a continuum-dominated band (García et al., 2013). However, observational studies imply that the X-ray source is typically radially extended (Krawczynski et al., 2022; Zhang et al., 2022; You et al., 2023; Wang & Zhang, 2024) and/or vertically extended (Kylafis et al., 2008; Ma et al., 2021; You et al., 2021; Méndez et al., 2022). Wilkins et al. (2016) provided a model that successfully reproduced the low-frequency hard lags and high-frequency soft lags with a corona both radially and vertically extended by assuming an energy descent in the corona and the bulk effect of the corona. Though this model explains the observed lags patterns with reverberation lags, spectral fitting indicates multiple emission components within the low-energy band (De Marco et al., 2021; Wang et al., 2021; You et al., 2021), suggesting that inter-band lags may result from these multiple components. Thus, a quantitative simulation for frequency-dependent lag spectra that considers an extended corona and self-consistent radiative transfer is necessary. Furthermore, a comprehensive model is essential to explore the connection between the inner accretion flow and the evolution of the lags.
@@ 2.1 Geometry
Figure 1: Cross section of the hyperbolic corona floating above the black hole. The upper and lower boundaries are controlled by the height parameters, $h_{t},h_{b}$. The outer surface is a hyperboloid with semi-major axis $a$ and semi-minor axis $b$
We consider a geometrically thin and optically thick disc that extends from the inner radius $r_{\text{in}}$ to the outer radius $r_{\text{out}}$. We mathematically model the corona as a hyperbolic surface with lower and upper surfaces (see Fig. 1) , which account for the extended geometry. This geometry is most readily interpreted as the base of the jet, although a more realistic jet base geometry would be parabolic as predicted in the simulation (Sridhar et al., 2025). Whether this geometry is vertically or radially extended with respect to the disc depends on the parameters used. The hyperboloid is parameterized by the following equations:
@@
where $a$ and $b$ are the semi-major and semi-minor axis of the hyperboloid, respectively. $h_{t}$ and $h_{b}$ are the top and bottom boundaries in $z$-direction. $\{x,y,z\}$ are the Cartesian coordinates for the black hole frame. We mainly consider that the corona is wider than it is tall, with the parameters $a=10R_{\rm g},b=2R_{\rm g}$, where $R_{\rm g}$ represents gravitational radii. This geometry would ensure a great number of thermal photons from the disc to enter the corona, thereby resulting in the Comptonization emission becoming predominant. An alternative spherical corona case will be discussed in Sect. 5. Some key parameters are shown in Table 1, with their definition and default values.
@@ 3 Monte Carlo Simulation
The code traces one single photon at a time. The trajectories of the photons are determined geometrically within the Newtonian framework for simplicity, without considering relativistic effects in the present work, e.g., time dilation and the light bending. Time dilation can be described by the gravitational redshift factor (g-factor), defined as the ratio between the frequency of observed and emitted photons. A g-factor of 1 indicates that there is no time dilation. Theoretical calculations in Kerr spacetime demonstrate that g-factor remains around 0.8-1.1 beyond $20R_{g}$, indicating only mild time dilation relative to flat spacetime (Dovciak, 2004). The light bending effect mainly affects the thermal emission and the reverberation process. Due to this effect, the radiation from the disc may return to illuminate itself, a phenomenon known as the returning effect. This process modifies the illumination profile to the disc, thereby altering both thermal and reflection spectra. However, the fraction of photons returning to the disc is significant only at very small radii $(<3R_{g})$, while this fraction is only $3\%$ at $20R_{g}$ for an extremely spinning black hole (Dauser et al., 2022). Spectral fitting further demonstrates minimal differences in parameter estimation when including or excluding the returning effect (Huang et al., 2025). Besides, light bending redistributes coronal illumination, concentrating photons toward the inner region of the disc. Our simulation with a $10R_{g}$-height lamp-post corona indicates a $\sim 14\%$ increase in illuminated photons within $30R_{g}$ and $\sim 8.7\%$ within $50R_{g}$, compared to the Newtonian case. The caveat is that such an altered illumination profile could influence the traveling distance of the photons and the reflection spectra, particularly when considering the ionization gradient and density gradient. This may modify inter-band time lags (Chainakun & Young, 2015). Moreover, some photons may be gravitationally bound around the black hole, undergoing multiple interactions between the corona and disc. This may form a thermal equilibrium — heating the inner disc while cooling the inner corona — thereby affecting the thermal emission and the Comptonization process. This effect is further complicated and is beyond the scope of this work. Thus, most relativistic effects are negligible; however, some effects remain uncertain regarding their significance. Future models will improve by incorporating these relativistic effects.
@@
A fraction of photons from the disc will strike the corona. These photons may undergo Comptonization, which would contribute to the Compton component. For simplicity, this work does not consider the effects of outflow or inflow velocity, nor the precession of the corona. Notably, the outflow/inflow velocity of the corona would change the seed photon luminosity due to the beaming effects, potentially influencing the simulated results(Malzac et al., 2001; Wilkins et al., 2016; You et al., 2021). Additionally, we assume a uniform plasma density and temperature within the corona, for simplicity. A density gradient leads to a gradient on optical depth. So the traveling distance of photons inside the corona would be changed compared to the case with uniform density, resulting in a different spatial lag. Similarly, a temperature gradient results in the energy of escaped photons from the corona being radially or vertically dependent, which would contribute to lags, as demonstrated in Wilkins et al. (2016), who assumed a photon index gradient in the corona. The optical depth of the corona is set in the range of 1-3, which ensures the spectrum is hard.
The simulation of Comptonization follows the standard methods established by Pozdnyakov et al. (1983) and Gorecki & Wilczewski (1984), which are parameterized by several factors related to the properties of the corona, such as optical depth and electron density, to determine whether the photons will be scattered.
@@ 3.1 Implementation of disc variability
In this study, we propose that the rapid variability originates from fluctuations in the accretion rate of the disc. To model the fluctuating accretion flow, we utilize the fluctuation model proposed by Ingram & Done (2011), which addresses local fluctuations and their propagation. The fluctuations propagate inward from the outer radius $r_{\text{out}}$ to the inner radius $r_{\text{in}}$ through the accretion flow, merging with local fluctuations along the way. For simplicity, we divide the accretion flow into $N_{\rm r}$ radial bins, each centered at $r_{n}$ in a logarithmic scale. The power spectrum of the local fluctuation at radius $r_{n}$ is represented by a zero-centered Lorentzian function:
@@
where $f$ is frequency, $f_{\mathrm{visc}}(r_{n})=f_{\max}(r_{n}/r_{i})^{-3/2}$ is the viscous frequency at radius $r_{n}$, in which $r_{i}$ is the inner edge of the disc (Shakura & Sunyaev, 1973). $f_{\max}$ is a free model parameter that sets the maximum viscous frequency in the disc.
@@ 3.2 Implementation of disc reflection
Partial Comptonized photons from the corona are reflected by the disc, resulting in both emission line and continuum emission. In the works of You et al. (2018, 2020), the reflection for a neutral disc was simulated to study the QPO variation. However, previous studies on BHXRBs have suggested the presence of an ionized disc (You et al., 2021). It has been shown that the flux in the soft band, for instance, 0.1-1 keV, is highly dependent on the ionization state of the disc (García et al., 2013). Therefore, simulating the reflection of an ionized disc is essential. This involves modeling various processes, such as absorption, excitation, and photoionization, which can be quite time-consuming. To address this, we adopt the table model Xillver (García et al., 2013) to simulate disc reflection.
[…]
@@
4 Result
We aim to simulate the inter-band lags across Fourier frequencies in the BHXRB. This will be achieved through Monte Carlo simulations of the rapid variability within the disc-corona system, taking into account both propagating fluctuations and reverberation mapping. Our goal is to quantitatively test whether these propagating fluctuations and reverberation mapping can explain the observed hard lag at low frequencies and the soft lag at high frequencies. Additionally, we will compare the results of our simulations with the observed evolution of frequency-dependent lags, which will help us constrain the geometry of the accretion flow and its evolution during outbursts. In this work, we will conduct simulations and comparisons specifically for MAXI J1820+070, which has provided valuable observations of fast variability (Kara et al., 2019; De Marco et al., 2021; Kawamura et al., 2023).
[…]
@@ 4 Result
The time-averaged energy spectrum of the three components is presented in Fig. 3 for demonstration purposes. The spectrum exhibits a cutoff power-law component with photon index $\sim 2$, a thermal component with the peak temperature $\sim 0.1$ keV, and a high-ionized reflection component. These features are consistent with the spectral fitting results, e.g., You et al. (2021) and Gao et al. (2023). We focus on an energy band of interest ranging from 0.1 to 0.3 keV, where all components are comparable, and a reference band from 2 to 5 keV, where the Compton component is prominent. The lag spectra between these two bands can be calculated using the Python package Stingray (Huppenkothen et al., 2019), as shown in Fig. 4. The errors arise from the incoherent components between the two bands and the stochastic nature of the variability (Bendat & Piersol, 2011; Uttley et al., 2014). The solid line in the plot represents a linear interpolation to guide the eye. We observe a low-frequency hard lag and a high-frequency soft lag in the spectrum, which are consistent with observations(Kara et al., 2019; Yu et al., 2023; De Marco et al., 2021).
@@
A trend towards shorter timescales for the lags is observed in MAXI J1820+070 as the accretion rate increases (Kara et al., 2019). Moreover, De Marco et al. (2021) note that the movement of the zero-crossing points from low to high frequencies indicates an evolution of the physical properties of the inner accretion flow (Alston et al., 2020; Wang et al., 2022a; Kalemci et al., 2022). We attribute these physical properties to the geometry of the corona and the inner viscosity. Our simulations captured the evolution of the corona’s geometrical size and viscosity frequency, with the corona’s upper bound decreasing from $35~{}R_{\rm g}$ to $16~{}R_{\rm g}$, accompanied by an increase in the maximum viscous frequency from 2 Hz to 30 Hz. Four evolving states are simulated, labeled state-1 to state-4, where state-1 is characterized by a maximum viscous frequency of 2 Hz and an upper bound for the corona of $35~{}R_{\rm g}$. The lag spectra between 0.1-0.3 keV and 2-5 keV for the four states are presented in Fig. 5.
@@ 5 Discussion
In typical timing spectral analysis, the energy-dependent lag spectrum is a good probe to understand the physical processes behind the lags at different energies in a specific frequency band (Arévalo et al., 2006; Cassatella et al., 2012; De Marco et al., 2015). The energy-dependent lag spectrum of MAXI J1820+070 shows that in the low frequency of 0.1-1 Hz, the lags increase as the energies increase in the energy band of 0.5-1 keV (Kara et al., 2019). This may originate from the propagation of fluctuations from a low energy area to high energy area, causing propagating lags. We tried to reproduce the energy-dependent lag spectrum from our simulation. We found that the simulated lag matches the the trend of the observed spectrum in the energy band below 1 keV, however, it breaks to a constant in the energy band over 1 keV. It’s necessary to note that the disk component vanishes over 1 keV, as shown in Fig. 3, and propagation of fluctuation in the corona is not considered in this work. Thus, after the disk component vanishes, the fluctuation can not propagate to the higher energy area, which may explain why the lags do not increase over 1 keV. We will implement the propagation of fluctuation in the corona in future work.